Insider Loans of nationally chartered banks across the country is approximately normally distributed with a mean of $500,000 and a standard deviation of $142,000. Find the probability that a nationally chartered bank selected at random has insider loans totaling at least $900,000.
Answer #1

Solution :

Given ,

mean = = 500000

standard deviation = = 142000

P(x >900000 ) = 1 - P(x< 900000)

= 1 - P[(x -) / < (900000-500000) / 142000]

= 1 - P(z < 2.82)

Using z table

= 1 - 0.9976

= 0.0024

probability= 0.0024